Could someone assist me with this question?
In the 19th century, J. B. A. Dumas devised a method for finding the molar mass of a volatile liquid from the volume, temperature, pressure, and mass of its vapor. (next column). He placed a sample of such a liquid in a flask that was closed with a stopper fitted with a narrow tube, immersed the flask in a hot water bath to vaporize the liquid, and then cooled the flask. Find the molar mass of a volatile liquid from the following:
Mass of empty flask = 65.347 g
Mass of flask filled with water at 25°C = 327.4 g
Density of water at 25°C = 0.997 g/mL
Mass of flask plus condensed unknown liquid = 65.739 g
Barometric pressure = 107.7 kPa
Temperature of water bath = 99.8°C
Answer #1

24.4-65.341 262.o 53 262.5 3 =-26289: mL 0.26284 L ( IL=1000 m L) ニ = 65.739-65.347 = 0.39 2 pv= nRT. V= o:26284L R= 0.0821 L